# MMAR: Towards Lossless Multi-Modal Auto-Regressive Probabilistic Modeling

### v2

@@ 4.3 Comparison with Other Systems
Gen. Only DALL-E 12B 250M 27.50
LDM 1.4B 400M 12.64
DALL-E2 6.5B 650M 10.39
Imagegen 3B 5000M+ 6.61
Und. and Gen. w/ pre-trained Diff. CoDI - 400M 11.26
SEED-X 17B - 12.68
DreamLLM 7B - 8.76
Joint Prob. Models Show-o 1.3B 35M 9.24
Chanmeleon 7B - 29.6
Transfusion  [^2] Both the Transfusion and Chanmeleon results are referenced from Table 3 in the paper ‘Transfusion: Predict the Next Token and Diffuse Images with One Multi-Modal Model.’ 7B - 16.8
@@
MMAR-7B 7B 145.2M 17.1
However, to bolster the model’s comprehension capabilities, exposure to diverse data is crucial rather than simply reiterating the same data. Consequently, we restrict our training to just 3 epochs for the second stage high-quality dataset. Despite these constraints, our model maintains competitive performance in generation, further substantiating the efficacy of our image-text joint modeling approach. The example images generated by our model are shown in the Appendix A.6.
Scaling up with Model Size. From Table 1 and 2, we see that MMAR can scale up from 0.5B to at least 7B model size, with significant improvement of visual understanding and generation capability.
4.4 Ablation Study
We begin by evaluating the impact of our chosen diffusion parameterization method. Table 3 demonstrates that switching to the more common $\epsilon$-prediction leads to a significant decrease in both visual understanding and generation quality, confirming the effectiveness of our optimal diffusion parameterization. To evaluate the efficacy of various image-text joint modeling techniques, we devise two distinct versions based on the MMAR framework: one employing LDM-VQ-16 tokenizer (Rombach et al., 2022) for discrete token modeling, and the other utilizing Transformer for diffusion modeling (refer to Transfusion). Comprehensive implementation details are provided in the Appendix A.1. Our test results are presented in Table 3. The Transformer-based diffusion modeling version considerably underperforms the other two approaches in both understanding and generation aspects. This is attributed to the substantial loss of image information when jointly modeling image-text and operating with limited training epochs. In contrast, our method consistently delivers superior results.
Table 3: Ablation study on MMAR
[…]
MMAR-0.5B(Full method) 48.45 882.1 70.74 55.70 18.49 34.56 36.6
w/ $\epsilon$-pred. 45.53 880.7 71.14 53.72 17.98 32.21 61.53
w/ VQ 37.54 618.2 66.98 44.93 14.45 29.70 66.26
transfusion-like 29.47 594.3 66.90 42.40 13.90 28.26 95.38
@@ 4.5 Analysis
To delve deeper into the disparities between v-prediction and $\epsilon$-prediction for diffusion MLP, we independently collect the statistics of the MSE loss of $v^{(t)}$ values at various time steps $t$ throughout the training process for both methods, as illustrated in fig. 3 (A). Furthermore, to more effectively discern the loss discrepancies between the two techniques, we subtract the v-prediction curve from the $\epsilon$-prediction curve, yielding the yellow curve. For reference, we plot the theoretically predicted numerical error of the $\epsilon$-prediction model (see Appendix A.3) as the red curve. The graph reveals that the loss of $\epsilon$-prediction model is consistently higher than v-prediction model. Especially, when $t>900$, curve “$\epsilon-v$” exhibits a significant spike towards infinity. This aligns with the behavior of the theoretical numerical error, confirming the non-negligible numerical error effect in low-precision training. The gap between yellow and red curve indicates that apart from its direct effect, numerical error also introduces optimization difficulty, hindering the loss convergence.
Impact of CFG Scaling
We select models from the second and fourth epochs of the first stage as starting points for the second stage, train them for 3 epochs, and then test the MSCOCO FID-30K under varying CFG intensities. As shown in fig. 3 (B), our method achieves better FID scores as the CFG scale increases from 1 to 10. It is worth noting that most probabilistic generative models typically have a CFG scale between 1.5 and 5. Additionally, it is observed that a longer training duration in the first stage consistently results in better generation outcomes at all CFG scales.
Scaling up with More Training Data
As illustrated in fig. 3 (C), we evaluate the average performance of MMAR-0.5B on 18 visual understanding benchmarks, following SFT training applied to the checkpoints generated during its training process. The blue background in the figure denotes the Image Expert Pretraining stage, while the green background signifies the Image Expert Fine-tuning stage. The curve reveals that, with an increasing number of training steps, i.e. more training data, the visual understanding performance of MMAR-0.5B consistently improves without reaching saturation. This finding highlights the exceptional scale-up capability of our MMAR model.
@@
Figure 3: The impact of v-prediction, CFG scaling, and training steps.
4.6 Limitation
Our method still requires further optimization in terms of image generation speed. Although the diffusion mlp is not involved in image understanding tasks, when applied to image generation tasks, we are compelled to use a 256-step token generation followed by a 100-step diffusion denoising process to ensure the quality of the generated images. This results in a generation time of nearly three minutes for a single 256x256 image using MMAR-7B. While it is possible to generate multiple images simultaneously by increasing the batch size, this does not fundamentally resolve the issue of prolonged generation time. We plan to address this in future work.
5 Conclusion
This paper proposes MMAR, a novel multimodal auto-regressive probabilistic modeling framework based on continuous image representations. It employs a standalone diffusion mlp at the image token level on top of pre-trained LLMs to facilitate theoretical lossless joint image-text probabilistic modeling. In practice, the low precision training of LLMs poses an non-negligible numerical error term to diffusion loss, causing optimization difficulty. This was addressed by deriving an optimal diffusion model parameterization. To balance the understanding and generation ability, a two-stage training strategy is introduced. During inference time, MMAR can tolerant extremely large CFG scale to generate high quality images. MMAR significantly demonstrates scaling-up laws with more data and larger model size. Extensive evaluations are conducted on 18 image understanding benchmarks, revealing that MMAR is the first joint image-text modeling framework that approaches comparable performance with traditional MLLMs that employ pretrained CLIP vision encoder, marking a significant step toward lossless joint probabilistic modeling of images and text.
Appendix A Appendix
A.1 Additional Implementation Details
VQ
@@ A.1 Additional Implementation Details
Based on the MMAR-0.5B framework, we replace the Image Tokenizer from KL-16 to VQ-16. The image codes extracted using VQ-16 are then passed through a projector to increase the channel size to match the LLM’s hidden size. Subsequently, we add a decoding Linear layer, which takes the hidden states of the LLM’s output image portion as input and maps them to the image codebook. The Cross Entropy loss is then calculated between these mapped values and the actual VQ codes.
Transfusion
Following the theoretical ideas presented in the Transfusion paper, we adopt a simple linear mapping. After extracting image tokens using KL-16, if the task is image first, we add noise within a 500-time step to the image tokens. Otherwise, we add noise within a 1000-step. After the linear mapping, we add the image token a learnable time embedding corresponding to the time step as input to the LLM. We also maintain the bidirectional attention mechanism.
After passing through the LLM, we first map the output back to the original token channel count using a linear layer and compute the MSE loss for the predicted noise. During generation, we treat LLM as a denoised model, with the condition being the concatenation of the text and the image tokens to be generated. We adopt the learning approach of Transfusion but conduct experiments based on our training tasks and stage divisions.
Projector
In order to accomplish channel alignment, we introduced two Projectors designed to scale the channels. Both Projectors consist of a simple linear layer and multiple blocks composed of activation layers and linear layers. The PostProjector comprises one block, whereas the VisProjector contains two blocks.
A.2 Minimizing the Numerical Error in Diffusion Models
@@
Now we consider the general form of DDIM sampling step (Song et al., 2021):
@@
$x^{(t-1)}$ $=\cos\phi_{t-1}\frac{\sin\psi_{t}x^{(t)}-\sin\phi_{t}u_{\theta}^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})}-\sin\phi_{t-1}\frac{\cos\psi_{t}x^{(t)}-\cos\phi_{t}u_{\theta}^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})}$
@@
Notably, not all these four factors are useful to achieve the goal of minimizing the numerical error. For example, tuning down the step size only decreases the numerical error of each step. As a result, the total step number of DDIM sampling is increased proportionally, which cancels out the effect of error reduction of each single step. The factor $u_{\theta}^{(t)}/r_{t}$ is not adjustable since it constantly has a unit standard deviation. This can be verified by the following calculation:
$\mathbb{E}[(u^{(t)}/r_{t})^{2}]=\mathbb{E}[(\cos\psi_{t}x+\sin\psi_{t}\epsilon)^{2}]=\cos^{2}\psi_{t}\mathbb{E}[x^{2}]+\sin^{2}\psi_{t}.$ (19)
In common practice, image tokens $x$ are normalized into unit standard deviation. Therefore, $\mathbb{E}[(u^{(t)}/r_{t})^{2}]=\cos^{2}\psi_{t}+\sin^{2}\psi_{t}=1$.
@@
We notice that this parameterization is coincidentally the “v-prediction” parameterization (Salimans & Ho, 2022). Note that, however, “v-prediction” is initially proposed for the efficient distillation of diffusion models, rather than reducing the numerical error of diffusion models. To the best of our knowledge, our work is the first to derive “v-prediction” parameterization from the first principle of minimizing the numerical error in diffusion models.
A.3 Deriving Theoretical Numerical Error for $\epsilon$-Prediction Models
@@
$v^{(t)}_{\theta}=\cos\phi_{t}\epsilon_{\theta}-\sin\phi_{t}\hat{x}_{\theta}(x^{(t)})=\cos\phi_{t}\epsilon_{\theta}-\sin\phi_{t}\frac{x^{(t)}-\sin\phi_{t}\epsilon_{\theta}}{\cos\phi_{t}}=\frac{\epsilon_{\theta}}{\cos\phi_{t}}-\tan\phi_{t}x^{(t)}.$ (24)
@@
$\mathbb{E}[(v^{(t)}-\tilde{v}^{(t)}_{\theta})^{2}]=\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta}-e^{(t)})^{2}]=\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta})^{2}]-2\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta})e^{(t)}]+\mathbb{E}[(e^{(t)})^{2}].$ (26)
@@
A.4 CFG With $v$-prediction

### v3

@@ 4.4 Ablation Study
Gen. Only DALL-E [45] 12B 250M 27.50
LDM [47] 1.4B 400M 12.64
DALL-E2 [46] 6.5B 650M 10.39
Imagegen [48] 3B 5000M+ 6.61
Und. and Gen. w/ pre-trained Diff. CoDI [56] - 400M 11.26
SEED-X [16] 17B - 12.68
DreamLLM [9] 7B - 8.76
Joint Prob. Models Show-o [62] 1.3B 35M 9.24
Chanmeleon [57] 7B - 29.6
Transfusion [71]  [^2] Both the Transfusion and Chanmeleon results are referenced from Table 3 in the paper ‘Transfusion: Predict the Next Token and Diffuse Images with One Multi-Modal Model.’ 7B - 16.8
@@
MMAR-7B 7B 145.2M 22.9
@@ 4.5 Analysis
To delve deeper into the disparities between v-prediction and $\epsilon$-prediction for diffusion MLP, we independently collect the statistics of the MSE loss of $v^{(t)}$ values at various timesteps $t$ throughout the training process for both methods, as illustrated in Fig. 5 (A). Furthermore, to effectively discern the loss discrepancies between the two techniques, we plot the difference between v-prediction and the $\epsilon$-prediction curve, yielding the yellow curve. For reference, we plot the theoretical numerical error of the $\epsilon$-prediction model (details in Appendix 6.3) as the red curve. The graph reveals that the loss of $\epsilon$-prediction model is consistently higher than v-prediction model. Especially, when $t>900$, curve “$\epsilon-v$” exhibits a significant spike towards infinity. This aligns with the behavior of the theoretical numerical error, confirming the non-negligible numerical error effect in low-precision training. The gap between yellow and red curve indicates that apart from its direct effect, numerical error also introduces optimization difficulty, hindering the loss convergence.
Table 4: Comparison on MJHQ-30K Dataset and GenEval.
@@
Gen. Only PixArt [3] 6.14 0.48
SDXL [43] 9.55 0.55
SD3 [14] - 0.68
Und. and Gen. w/ pre-trained Diff. CoDI [56] - 0.31
SEED-X [16] - 0.55
Joint Prob. Models Show-o [62] 15.18 0.53
VILA-U [61] 12.81 -
Transfusion [71] - 0.63
EMU3 [60] - 0.66
MMAR-7B 15.6 0.51
Scaling Law
We conduct scaling up experiments by varying model size (N) and training token count (D), approximating FLOPs as 6ND. Fig. 5 (B), (C) displays the performance on image understanding and generation benchmarks. The scaling laws are evident in both tasks, as shown by the dashed lines.
[…]
Figure 5: The impact of v-prediction and the scaling laws on the tasks of image understanding and generation.
5 Discussion and Conclusion
MMAR is the first to achieve the unified autoregressive probabilistic modeling for continuous image and discrete text representation. It addresses numerical instability and balances image comprehension and generation, thereby ensuring practicality and scalability. While masked autoregressive modeling serves as one of the design options for MMAR, our framework can be adaptable to future breakthroughs in continuous image autoregressive modeling. MMAR uses continuous image tokens, boosting efficiency sixteenfold over image discretization, matching or exceeding EMU-3’s image comprehension capabilities with fewer image tokens (4096 tokens/image to 256 tokens/image). This advantage grows with advanced continuous image tokenizers. And it would definitely benefit more when it comes to modeling of interleaved multi-image data and long videos. MMAR is currently in early exploration, being experimented on relatively limited resolution, data scale, model size, and did not utilize more advanced continuous image tokenizers. Despite this, MMAR shows promise through its scalable performance and strong comprehension/generation abilities. Future works include unlocking these limitations and developing more advanced continuous image AR implementation other than MAR.
Acknowledgments
This work is supported by the National Natural Science Foundation of China (NSFC) under Grants 62225207, 62436008 and 62306295. Additionally, we thank Ziheng Zhou for his helpful discussion.
Supplementary Material
6 Appendix
6.1 Additional Implementation Details
Table 5: Hyper-parameter settings for MMAR models
MMAR-0.5B MMAR-7B
Module Param. settings settings
$N_{diff}$ 8 12
$d_{mlp}$ 1024 2048
Diffusion MLP $r_{mlp}$ 4 4
Res. 256 256
Image Tonkenizer $d$ 16 16
$N_{ViT}$ 16 16
EmbeddingViT $d_{ViT}$ 1024 1024
$r$ 512 1280
LLM’s PLoRA $\alpha$ 128 128
Diffusion MLP.
Inspired by MAR, we employ a simple MLP architecture to predict $v^{(t)}$, whose detailed architecture is shown in Fig. 6. It consists of a main network and its input/output linear projection layers. The input projection converts the $d$-dimensional noisy latent $x^{(t)}$ into the mlp hidden dimension $d_{mlp}$, while the output projection converts the $d_{mlp}$-dimensional main network output back to $d$ dimensions. The main network is a stack of $N_{diff}$ residual blocks, each comprising an AdaLN [42], followed by a two-layer MLP activated by SiLU. The expand ratio of each MLP is denoted by $r_{mlp}$, which means its first linear layer projects from $d_{mlp}$ to the dimension of $d_{mlp}\times r_{mlp}$. The condition vector $z$ is added to the diffusion time embedding and is incorporated through AdaLN. In practice, $r_{mlp}$ is set to 4. For MMAR-0.5B, $N_{diff}=8$ and $d_{mlp}=1024$, while for MMAR-7B, $N_{diff}=12$ and $d_{mlp}=2048$.
Image Tokenizer.
MMAR employs the publicly available KL-16 image tokenizer from LDM [47]. This tokenizer processes an 256 $\times$ 256 image into 256 image tokens, each with $d=16$ channels, and it is kept frozen during training.
EmbeddingViT.
The EmbeddingViT module is implemented with a ViT[11] encoder with $N_{ViT}=16$ layers and $d_{ViT}=1024$ hidden state channels, processing image tokens into visual embeddings with stronger context awareness. We integrate a learnable position embedding for each image token in EmbeddingViT, corresponding to its 2D position on the image.
LLM.
We initialize our LLM using parameters from the open-source QWen2 series models [63]. To preserve text modeling capabilities, we keep the LLM parameters fixed during training and add PLoRA [10] as the image expert, where only the image tokens in the input pass through the introduced LoRA [22] adapters. The PLoRA is applied for each linear layer within the original LLM. Considering the time cost, our ablation study employs QWen2-0.5B-Instruct, with $r=512$ and $\alpha=128$ for PLoRA. Furthermore, we use QWen2-7B-Instruct to explore our method’s scale-up capability with $r=1280$ and $\alpha=128$ for PLoRA. The corresponding MMAR models are denoted by MMAR-0.5B and MMAR-7B, respectively.
Figure 6: Details of Diffusion MLP architecture.
Projector.
We use two MLP-based projectors [68, 69, 67] to connect different modules’ input and output. Specifically, the PostProjector connects the Image Tokenizer output to the EmbeddingViT’s input with a 2-layer MLP, whereas the VisProjector connects EmbeddingViT output to the LLM’s input with a 3-layer MLP.
Training Task Proportion.
MMAR models $P(I|T)$, $P(T|I)$ and $P(I)$ through the conditional image generation task, the image understanding task and the unconditional image generation task, respectively. Through the first and the second stage training, we set the proportion of these three tasks as 50%, 45%, and 5%, respectively. During the optional third training stage, only image generation tasks are applied, and the proportion of the conditional and unconditional image generation tasks are set to 90% and 10%, respectively.
Mixed-Precision Inference.
To enhance numerical stability during image generation, particularly when using small step sizes in the DDPM sampling process, we perform the model’s forward pass in bfloat16 (matching the training precision) but cast the output to float32 before DDPM sampling. This mitigates potential rounding error without significant computational overhead, improving sampling accuracy efficiently.
Ablation-VQ.
@@ 6.1 Additional Implementation Details
Based on the MMAR-0.5B framework, we replace the Image Tokenizer from LDM-KL-16 to LDM-VQ-16. The image codes extracted using LDM-VQ-16 are then passed through a projector to increase the channel size to match the LLM’s hidden size. Subsequently, we add a decoding Linear layer, which takes the hidden states of the LLM’s output image portion as input and maps them to the image codebook. The Cross Entropy loss is then calculated between these mapped values and the ground-truth VQ codes.
Transfusion Reimplementation.
Our re-implementation version of Transfusion shares most of the MMAR-0.5B’s model architecture except for the processing of the model input and output. Following the Transfusion paper, we adopt a input linear projection to convert the output of LDM-KL-16 image tokenizer into the LLM input representation, and use an output linear projection to convert the LLM output into the predicted noise. Noise of different levels are added to the image token input according to different pretraining tasks. For the image-to-text task, the diffusion time step is uniformly sampled within $t\in[0,500]$, while for the text-to-image task, the diffusion time step is uniformly sampled within $t\in[0,1000]$. A learnable time embedding corresponding to the time step $t$ is added after the input linear projection. MSE loss is calculated between the predicted and ground-truth noise. During inference, we treat the LLM as a diffusion model, with the condition being the concatenation of the text and the noisy image tokens of the previous diffusion time step $t+1$.
6.2 Minimizing the Numerical Error in Diffusion Models
@@
Now we consider the general form of DDIM sampling step [51]:
@@
$x^{(t-1)}=$ $\cos\phi_{t-1}\frac{\sin\psi_{t}x^{(t)}-\sin\phi_{t}u_{\theta}^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})}$
$-\sin\phi_{t-1}\frac{\cos\psi_{t}x^{(t)}-\cos\phi_{t}u_{\theta}^{(t)}/r_{t}}{\sin(\psi_{t}-\phi_{t})}$
@@
Notably, not all these four factors are useful to achieve the goal of minimizing the numerical error. For example, tuning down the step size only decreases the numerical error of each step. As a result, the total step number of DDIM sampling is increased proportionally, which cancels out the effect of error reduction of each single step. The factor $u_{\theta}^{(t)}/r_{t}$, which represents the normalized model prediction, is largely decided by the optimization progress and the target data distribution. An approximation can be derived from a perfect prediction, that is $u_{\theta}^{(t)}/r_{t}=u^{(t)}/r_{t}$. In this way,
$\mathbb{E}[(u_{\theta}^{(t)}/r_{t})^{2}]$ $\approx\mathbb{E}[(\cos\psi_{t}x+\sin\psi_{t}\epsilon)^{2}]$
$=\cos^{2}\psi_{t}\mathbb{E}[x^{2}]+\sin^{2}\psi_{t}.$ (19)
In common practice, image tokens $x$ are normalized into unit standard deviation, leading to $\mathbb{E}[(u_{\theta}^{(t)}/r_{t})^{2}]\approx\cos^{2}\psi_{t}+\sin^{2}\psi_{t}=1$. This is a constant number, meaning that there is not much potential to reduce the total numerical error via tuning down $u_{\theta}^{(t)}/r_{t}$.
@@
We notice that this parameterization is coincidentally the “v-prediction” parameterization [49]. Note that, however, “v-prediction” is initially proposed for the efficient distillation of diffusion models, rather than reducing the numerical error of diffusion models. To the best of our knowledge, our work is the first to derive “v-prediction” parameterization from the first principle of minimizing the numerical error in diffusion models.
Table 6: Detailed visual understanding evaluation results.
MMAR-0.5B MMAR-0.5B MMAR-7B
Benchmark Chameleon-7B Transfusion* Show-o MMAR-0.5B w/ $\epsilon$-pred. w/ VQ MMAR-7B w/ Stg3
AI2D [24] 34.81 40.22 32.48 43.43 41.90 41.90 64.64 63.54
ChartQA [40] 3.84 9.56 11.32 10.20 10.36 9.36 13.64 12.52
DocVQA [41] 1.51 6.72 18.24 7.62 6.77 6.79 11.12 10.62
Hallu.Bench [18] 39.01 41.54 40.90 42.80 41.11 41.54 53.10 52.05
MathVista [39] 21.90 22.60 23.20 21.60 23.10 22.90 32.40 31.30
MMBenchCN [35] 10.14 27.23 0.52 43.99 38.83 31.87 70.53 66.75
MMBenchEN [35] 13.32 29.47 42.44 48.45 45.53 37.54 70.45 67.78
MMEP [15] 125.8 594.3 1182.7 882.1 880.7 618.2 1486.9 1421.9
MMEC [15] 33.9 206.1 225.0 256.8 232.1 273.2 268.9 303.6
MMMU [66] 24.00 29.33 26.44 29.33 25.33 29.67 41.33 47.33
MMStar [5] 20.47 28.13 32.00 32.13 31.07 28.07 41.87 40.87
MMVet [65] 7.34 13.90 20.87 18.49 17.98 14.45 30.64 29.17
OCRBench [36] 0.50 2.30 15.20 18.70 7.10 2.10 25.80 21.20
POPE [31] 30.86 66.90 84.50 70.74 71.14 66.98 84.02 83.21
RealWorldQA 27.06 36.99 27.97 38.30 35.16 36.60 53.59 52.68
ScienceQA [38] 44.83 45.92 41.82 47.54 45.21 45.35 74.39 73.20
SEEDBench [26] 34.61 42.40 51.61 55.70 53.72 44.93 68.63 66.59
TextVQA [50] 5.43 9.94 38.35 16.77 12.40 9.46 24.37 21.35
Average 18.34 28.26 33.06 34.56 32.21 29.70 48.25 47.18
6.3 Deriving Theoretical Numerical Error for $\epsilon$-Prediction Models
@@
$v^{(t)}_{\theta}$ $=\cos\phi_{t}\epsilon_{\theta}-\sin\phi_{t}\hat{x}_{\theta}(x^{(t)})$
$=\cos\phi_{t}\epsilon_{\theta}-\sin\phi_{t}\frac{x^{(t)}-\sin\phi_{t}\epsilon_{\theta}}{\cos\phi_{t}}$
$=\frac{\epsilon_{\theta}}{\cos\phi_{t}}-\tan\phi_{t}x^{(t)}.$ (24)
@@
$\mathbb{E}[(v^{(t)}-\tilde{v}^{(t)}_{\theta})^{2}]=$ $\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta}-e^{(t)})^{2}]$
$=$ $\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta})^{2}]-2\mathbb{E}[(v^{(t)}-v^{(t)}_{\theta})e^{(t)}]$
$+\mathbb{E}[(e^{(t)})^{2}].$ (26)
@@
Figure 8: The impact of CFG scale on image generation quality.
6.4 CFG With $v$-prediction